OpenStudy (anonymous):

Calculate y^{(k)}(0) for 0 <= k <= 5, where y=6 x^4+a x^3 + b x^2 + c x + d (with a,b,c,d the constants) y^{(0)}(0)= y^{(1)}(0)= y^{(2)}(0)= y^{(3)}(0)= y^{(4)}(0)= y^{(5)}(0)=

OpenStudy (anonymous):

I know the first one is just d but I can't figure out the rest.

OpenStudy (aum):

Is y^k the kth derivative?

OpenStudy (aum):

Is y^1 dy/dx ? y^2 d^2y / dy^2 ?

OpenStudy (aum):

or is it just an exponent? y^k = \(y^k\) ?

OpenStudy (anonymous):

yeah its just an exponent I think

OpenStudy (anonymous):

If the exponent's wrapped in parentheses, they tend to denote orders of derivatives...

OpenStudy (freckles):

I would rather do derivative junk Looking it as an exponent seems like really mean work

OpenStudy (aum):

Well they are all evaluated at x = 0

OpenStudy (freckles):

like punishment work

OpenStudy (freckles):

oh that's true lol

OpenStudy (freckles):

then y(0)=d (y(0))^2=d^2

OpenStudy (aum):

Is this question from a calc class?

OpenStudy (freckles):

but I would think it means kth derivative

OpenStudy (anonymous):

yeah it is

OpenStudy (aum):

Then it is the derivative.

OpenStudy (aum):

\[ y=6 x^4+a x^3 + b x^2 + c x + d \\ y^{(0)}(0) = d \\ y^{(1)} = 24x^3 + 3ax^2 + 2bx + c \\ y^{(1)}(0) = c \\ y^{(2)} = ? \\ y^{(2)}(0) = ? \\ .... \]

OpenStudy (anonymous):

2b right?

OpenStudy (anonymous):

Ok thanks I can get the rest